QUESTION 10 On Jetstar flights, passengers have to pay extra for food and drinks. Assume the only product that Jetstar sells on the plane is packets of potato chips. Each packet costs Jetstar $4.00 (hint: the marginal cost is constant at $4.00 per packet). Currently Jetstar charges $4.50 for a packet of chips Price ($/packet) 6.50 6.00 Quantity (packets) 30 42 5.50 5.00 4.50 4.00 56 72 92 105 Answer the following questions a. What is Jetstars marginal revenue when the price is $4.50 per packet? Answer to the nearest two decimal places. $ per packet b. At a price of $4.50, should Jetstar decrease or increase the packets of chips sold to maximise profit? Type D for Decrease or l for Increase c. What price should Jetstar charge to maximise profits? Answer to the nearest two decimal places. $ per packet.
Answer #1

When price is 5.00, revenue is 72*5 = 360. When price is 4.50, revenue is 414. Hence marginal revenue when price is 4.50 is (414 - 360)/(92 - 72) = $2.70

Since marginal cost is 4.00, the current price is not profit maximizing as marginal revenue is lower than marginal cost. Hence it should decrease output

Marginal revenue when price is 5.50 is (5.5*56 - 6*42)/(56-42) = 4 which is equal to marginal cost. Hence it should charge a price of $5.50